• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Carlos

    Log entry time 01:19:44 on December 1,2004

    Entry number 137576

    keyword=Compton vertical scan

    Vertical lock was not in, so 102B setting drifted and rates were of around 500Hz when cavity was locked. I asked MCC to do a vertical scan and put the lock in. Rates are back to 900Hz-1000Hz.